A study uses a random sample of size 15. The test statistic for testing Upper H 0 : mu equals 12 versus Upper H Subscript a Baseline : mu not equals 12 is t equals negative 2.5 . Find the approximate​ P-value. Round your answer to two decimal places. A. 0.013 B. 0.050 C. 0.990 D. 0.025 E. The​ P-value cannot be determined without the sample standard deviation.

(please explain the steps using TI84 calculator and the keys to enter in the calculator!!!)
Answer #1

Option D) is correct.

P value is 0.025...............................By using Excel command TDIST(-2.5,14,2)

Degree of freedom =n-1 = 15-1 = 14

P value procedure by using TI-84
Find the absolute value of t_calc
>>2nd DISTR
>>Scroll down to tcdf(
>> ENTER
>> Now enter: |t_calc|, 1000, df)
>> ENTER
>>Output is ½ of the P-value so
>> Multiply result by 2